OpenStudy (anonymous):

A is an acute angle with tanA=1/2. The value of cos2A is what?

OpenStudy (anonymous):

lol your cute face sry no answer;(

OpenStudy (anonymous):

use Trig identities....

OpenStudy (anonymous):

using Half-angle Formula \[\tan^{2}u=\frac{ 1-\cos2u }{ 1+\cos2u }\]

OpenStudy (anonymous):

i use u instead of A... anyway it's just the same... you can solve it now by substituting tan A=1/2 or tan u = 1/2

OpenStudy (anonymous):

lol